# The canonical lamination calibrated by a cohomology class

### v3

@@ 4. Construction of the canonical lamination
Let $F$ be a calibration in $\rho$. The set $S:=\{|F|=1\}$ need not be the support of a lamination $\lambda$; and even if it was, we would not be able to conclude that $F$ calibrates $\lambda$. For example, if $d\geq 3$, then one can exploit the possible nonintegrability of $\star F$ to produce counterexamples [bangert_cui_2017, §4]. More starkly, if $d=2$, then the main theorem of [daskalopoulos2020transverse] then implies that $S$ contains a geodesic lamination $\lambda$; on the other hand, the main theorem of [backus2024lipschitz] implies that any closed set containing $\operatorname{supp}\lambda$ can be realized as the set $\{|G|=1\}$ for some calibration $G$ in $\rho$. If $M$ is hyperbolic, then $\lambda$ has Hausdorff dimension $1$ [Birman1985], so “almost every” closed subset of $M$ is $\{|G|=1\}$ for some $G$.
@@
We now show that the leaves of the putative canonical lamination satisfy the necessary curvature bounds; this is a little subtle because the leaves are injectively immersed but not embedded. In the below lemmata, let $r_{*}$ be the minimum of the injectivity radius of $M$ and $\delta\|\mathrm{Riem}_{M}\|_{C^{0}}^{-1/2}$, where $\delta>0$ is a dimensional constant to be determined later. Let $\mathbf{S}^{d-1}$ be the round sphere of dimension $d-1$.
@@
$\mathrm{vol}(\mathbf{B}^{d-1})\leq\mathrm{vol}(N^{\prime})\leq 2\mathrm{vol}(\mathbf{S}^{d-1})r^{d-1}.$ (4.1)
@@
Let $N^{\prime}$ be a component of $N\cap B(x,r)$, so that $N^{\prime}$ is embedded (not just injectively immersed). By transversality, $N^{\prime}\cap\partial B(x,r)$ is diffeomorphic to a closed $d-2$-dimensional submanifold of $\mathbf{S}^{d-1}$. Since $H_{d-2}(\mathbf{S}^{d-1},\mathbf{R})=0$, there exists a relatively open set $V\subseteq\partial B(x,r)$ which is bounded by $N\cap\partial B(x,r)$. Because of how we chose $r_{*}$, we may use the $L^{\infty}$ Poincaré lemma, Theorem 2.1, to find a continuous $d-2$-form $A$ on a neighborhood of the closure of $B(x,r)$, such that $F=\mathop{}\!\mathrm{d}A$. Then
@@
Let $x\in N$ and let $r>0$ be small enough depending on $M$. Then each component $N^{\prime}$ of $N\cap B(x,r)$ is $F$-calibrated, and therefore a stable minimal hypersurface. By (4.1), $\mathrm{vol}(N^{\prime})\lesssim r^{d-1}$. So by [Schoen81, pg785, Corollary 1] (see also [colding2011course, Chapter 2, §§4-5]),
@@
We first observe that for each $x\in N$, $(\star F(x))^{\sharp}$ is the (unique) normal vector to $N$ at $x$ (and similarly for $N^{\prime}$), and so if $x\in N\cap N^{\prime}$ then $N,N^{\prime}$ have the same tangent space at $x$. So for each $x\in N\cap N^{\prime}$, there exists $r>0$ and normal coordinates $(\xi,\eta)\in\mathbf{R}^{d-1}\times\mathbf{R}$ on $B(x,r)$ based at $x$, such that for each pair of sheets $N_{*}\subseteq N\cap B(x,r)$, $N_{*}^{\prime}\subseteq N^{\prime}\cap B(x,r)$ which contain $x$, there exists a relatively open set $V\subseteq\{\eta=0\}$, an open set $U\subseteq B(x,r)$ containing $x$, and functions $u,u^{\prime}:V\to\mathbf{R}$ such that:
@@
    $v$ satisfies a linear elliptic PDE on $V$ [colding2011course, Proof of Theorem 7.3].
@@
We claim that $v$ is identically 00. If this is not true, the set $\{v=0\}=\{v=\mathop{}\!\mathrm{d}v=0\}$ is $d-3$-rectifiable [Hardt89, Lemma 1.9], but $\dim V=d-1$, so $\{v\neq 0\}$ is connected. So either $v\geq 0$ or $v\leq 0$, and $v$ has a zero; this contradicts the maximum principle.
@@
We must show that a limit of $F$-calibrated hypersurfaces is $F$-calibrated, and to make this precise we shall need the notion of a Vietoris limit superior of a sequence of closed sets, [kechris2012classical, §4.F]. If $(K_{n})$ is a sequence of closed subsets of $M$, then $\limsup_{n\to\infty}K_{n}$ is the set of all $x$ such that for every open set $U\ni x$, there exist infinitely many $n\in\mathbf{N}$ such that $U\cap K_{n}\neq\emptyset$; one easily checks that $\limsup_{n\to\infty}K_{n}$ is closed.
@@
By taking a subsequence, we may assume that there exist $x_{n}\in N_{n}$ such that $x_{n}\to x$. By Lemma 4.4, we may also assume that if $N_{n}\cap N_{m}$ is nonempty then $n=m$. Combining this with the curvature bound (4.2), we obtain the hypotheses of [BackusCML, Lemma 2.4]. The conclusion of that lemma is that for every $\delta>0$ there exists $r>0$ only depending on $M$, and normal coordinates $(\xi,\eta)\in\mathbf{R}^{d-1}\times\mathbf{R}$ on $B(x,r)$ based at $x$ such that for every $n$,
@@
where one can use [colding2011course, (7.21)] to show that $F$ has the form
@@
So by (4.3), we can first choose $\delta$ small enough depending on $M$, and then choose $r$ small enough depending on $\delta$, so that for every $n$ large enough depending on $r$, the equation $Pu_{n}=0$ is uniformly elliptic. In particular, by the interior Schauder estimate [gilbarg2015elliptic, Theorem 6.2], we may choose $\delta$ small enough, depending only on $M$, that there exists a connected, relatively open set $W\subseteq V$ such that:
@@
To prove this, let $N_{n}^{x}:=\{(\xi,u_{n}(x)):\xi\in W\}$. If $\mathrm{diam}(W)$ was chosen small enough (depending only on $M$), then we can use the $L^{\infty}$ Poincaré lemma, Theorem 2.1, to find a continuous $d-2$-form $A$ on $W$ such that $\mathop{}\!\mathrm{d}A=F$. Since $u_{n}\to u$ in $C^{2}(W)$, we can compute using Stokes’ theorem
@@ 5.1. Ergodic theory of $\lambda_{\rho}$
Let $M$ be a closed oriented Riemannian manifold of dimension $2\leq d\leq 7$. For each oriented lamination $\lambda$ in $M$, let $\mathcal{M}(\lambda)$ be the set of transverse probability measures to sublaminations of $\lambda$. This set inherits the vague topology on the space of Borel probability measures on $\operatorname{supp}\lambda$; in view of (2.3), this topology is the same as the topology on the space of measured laminations (see [BackusCML]) restricted to $\mathcal{M}(\lambda)$. It is clear that $\mathcal{M}(\lambda)$ is convex, and one may use the compactness of the space of Borel probability measures on the compact metrizable space $\operatorname{supp}\lambda$ [kechris2012classical, Theorem 17.23] to show that $\mathcal{M}(\lambda)$ is compact. By the Krein-Milman theorem, if $\mathcal{M}(\lambda)$ is nonempty, then so is its set of extreme points, $\mathcal{E}(\mathcal{M}(\lambda))$.
@@
The first claim is an easy modification of the proof of [einsiedler2010ergodic, Theorem 4.4], and the second is essentially the proof that every ultrafilter on a finite set is principal. To be more precise, let $S$ be a finite set of ergodic measures, and choose $c_{\mu}\in\mathbf{R}$ such that $\sum_{\mu\in S}c_{\mu}\mu=0$. The measures in $S$ are determined by their values on saturated sets, so if some coefficient $c_{\nu}$ is nonzero, then there exists a saturated set $E$ and a proper subset $T\subset S$ such that:
@@ 5.2. The Arnoux–Levitt lemma
Following an idea of Auer and Bangert [Auer01], we study $\mathcal{M}(\lambda_{\rho})$ using an ergodic-theoretic lemma of Arnoux and Levitt, [Arnoux1986, Proposition 3.1]. We need a more general version of the Arnoux–Levitt lemma, and the original proof is in French, so we include a proof here.
Let us identify transverse measures with positive transverse cocycles (cocycles which act on curves transverse to the lamination and are cooriented with the lamination); this is standard, and we refer to [daskalopoulos2020transverse, §7.2] for a justification of this identification.
@@
We follow [Arnoux1986, §3, Lemme] which is a similar result when $\lambda$ is a minimal component of a foliation.
@@
    $N$ is dense in $\operatorname{supp}\lambda\cap U$.
@@
    Suppose that $\kappa_{i}$ is not a foliation with a dense leaf. By the Morgan–Shelan decomposition (Theorem 2.12) and the ergodicity of $\mu_{1}^{i}$, $\kappa_{i}$ is exceptional, and if $i^{\prime}\neq i$ then $\operatorname{supp}\kappa_{i^{\prime}}$ avoids an open set $U_{i}$ containing $\operatorname{supp}\kappa_{i}$. Let $N_{i}$ be any leaf of $\kappa_{i}$.
@@
Suppose that there are $a_{j}\in\mathbf{R}$ such that $\sum_{j}a_{j}[\mu^{i}_{j}]=0$, let $\mu$ be the sum of $a_{j}\mu_{i}^{j}$ over $j$ such that $a_{j}>0$, and let $\nu$ be the sum of $-a_{j}\mu_{i}^{j}$ over $j$ such that $a_{j}<0$. Then $[\mu]-[\nu]=0$, so by Lemma 5.3 with $b=0$ and $U=M$, $\mu-\nu=0$, hence $\sum_{j}a_{j}\mu^{i}_{j}=0$. By ergodicity, $(\mu_{i}^{j})_{j}$ is linearly independent, so $a_{j}=0$, establishing (0–1).
@@
$[\mu_{i}^{1}]:H_{1}(M,\mathbf{Z})\to\mathbf{R}.$
Since $M$ is compact, $H_{1}(M,\mathbf{Z})$ is finitely generated and so we can slightly perturb the value of $[\mu_{i}^{1}]$ on the generators to obtain $t_{i}$ with the desired properties.
@@
We have already seen that every extreme point of $\mathcal{M}(\lambda)$ is an ergodic measure and so we just need to show the converse. By the Morgan–Shelan decomposition, Theorem 2.12, we may reduce to the case that $\lambda$ either has only closed leaves, or has no closed leaves. If $\lambda$ has only closed leaves, then every ergodic sublamination is a single leaf, and in particular cannot be written as a convex combination of any other ergodic sublaminations. Otherwise, $\lambda$ has no closed leaves, so by Lemma 5.4, $\mathcal{E}(\mathcal{M}(\lambda))$ is finite, say $\{\mu_{1},\dots,\mu_{m}\}$. For each $i\neq j$, there exists a saturated Borel set $E_{ij}$ such that $\mu_{i}(E_{ij})=1$ but $\mu_{j}(E_{ij})=0$. In particular, if $\mu$ is not an extreme point and we write $\mu=\sum_{i}d_{i}\mu_{i}$, then there exist $i\neq j$ such that $d_{i},d_{j}>0$. Then $d_{i}\leq\mu(E_{ij})\leq 1-d_{j}$, so $\mu$ is not ergodic. ∎
@@ 5.3. The dual flat $\rho^{*}$
    $[\lambda]$ has rational direction.
@@
If $\lambda$ is a closed hypersurface $N$, then $[\lambda]$ is a rescaling of $[N]$, and $[N]$ is the image of the class of $N$ in $H_{d-1}(M,\mathbf{Z})$.
Conversely, assume that $[\lambda]$ has rational direction. Since $\Gamma$ is finitely generated, we may rescale $M$ suitably so that $[\lambda]$ is a representation $\alpha:\Gamma\to\mathbf{Z}$. Since such representations are identified with homotopy classes of maps $M\to\mathbf{S}^{1}$, the Ruelle-Sullivan current $T_{\lambda}$ takes the form $\mathop{}\!\mathrm{d}u$ for some map $u:M\to\mathbf{S}^{1}$. Let $\tilde{u}\in BV_{\mathrm{loc}}(\tilde{M},\mathbf{R})$ be the universal cover of $u$.
@@
To see this, let $\tilde{y}\in\partial\{u>t\}\cap B(\tilde{x},r/2)$. Since $\partial\{\tilde{u}>t\}$ is smooth, its density $\theta$ (in the sense of rectifiable sets) at $\tilde{y}$ is the volume of the unit ball of $\mathbf{R}^{d-1}$. By the monotonicity formula for minimal hypersurfaces [Marques, Theorem 7.11], there exists $A\geq 0$ which only depends on $M$ such that for any $\rho>0$,
@@
The image of $T$ in $\mathbf{S}^{1}$ is a point, so for any $t,s\in T$, either $t=s$ or $|t-s|\geq 1$. We may assume that there is an infinite increasing sequence $(t_{n})$ in $T$. By the coarea formula [Giusti77, Theorem 1.23],
@@
So if $[\lambda]$ has rational direction, then every leaf of $\lambda$ is closed. Since $\lambda$ is ergodic, it follows that $\lambda$ is a single closed leaf. ∎
@@
By Lemma 5.6, the map $\Pi:\mathcal{M}(\lambda_{\rho})\to\partial B$ which takes a transverse probability measure to its homology class sends closed hypersurfaces to rational vertices, and measures on sublaminations which are not closed hypersurfaces to irrational vertices. By Corollary 1.3, $\Pi$ is surjective, so in order to bound the number of irrational vertices, we need to bound the number $m$ of measures on sublaminations which are not closed hypersurfaces. So by Lemma 5.4, $m\leq\max(1,b_{1}-1)$.
To complete the proof, we must show that $\rho^{*}$ has finitely many rational vertices. If not, there are infinitely many closed leaves $N_{n}$ of $\lambda_{\rho}$ with distinct homology classes $\alpha_{n}\in\mathcal{E}(\rho^{*})$. The infinite sequence $(\alpha_{n})$ is linearly dependent, so $M\setminus\bigcup_{n}N_{n}$ must be disconnected; therefore there can be no leaf of $\lambda_{\rho}$ which is dense in $M$. There is a measure in $\mathcal{M}(\lambda_{\rho})$ which assigns each $N_{n}$ positive weight, so by the Morgan–Shelan decomposition (Theorem 2.12) and pigeonholing, there is $n\neq m$ such that $N_{n},N_{m}$ are in the same parallel family, so $\alpha_{n}=\alpha_{m}$, a contradiction. ∎
@@
5.4. Perimeter-minimizing sets
In this section only, $M$ denotes a complete Riemannian manifold of bounded curvature; we do not assume that $d=7$ or $M$ is closed. A Borel set $U\subseteq M$ is perimeter-minimizing if $1_{U}$ is a function of least gradient. In the proof of Theorem 1.6(2), we shall need an estimate on perimeter-minimizing sets, which we now prove. See [Giusti77, Chapter 5] for the proof when $M$ is an open subset of euclidean space.
Lemma 5.7.
$\mathrm{vol}(U\cap B(x,r))\geq cr^{d}.$ (5.1)
Proof.
where $c_{d}>0$ is the euclidean isoperimetric constant. We can reason as in the proof of [Giusti77, Proposition 5.14] to see that for almost every $0<\rho<r$,
Let $f(\rho):=\mathrm{vol}(U\cap B(x,\rho))^{1/d}$, so that
Therefore $f(\rho)\geq\rho/(4dc_{d})$, as desired. ∎
Lemma 5.8.
Proof.

### v4

@@ 4. Construction of the canonical lamination
Let $F$ be a calibration in $\rho$. The set $S:=\{|F|=1\}$ need not be the support of a lamination $\lambda$; and even if it was, we would not be able to conclude that $F$ calibrates $\lambda$. For example, if $d\geq 3$, then one can exploit the possible nonintegrability of $\star F$ to produce counterexamples [8, §4]. More starkly, if $d=2$, then the main theorem of [17] then implies that $S$ contains a geodesic lamination $\lambda$; on the other hand, the main theorem of [5] implies that any closed set containing $\operatorname{supp}\lambda$ can be realized as the set $\{|G|=1\}$ for some calibration $G$ in $\rho$. If $M$ is hyperbolic, then $\lambda$ has Hausdorff dimension $1$ [9], so “almost every” closed subset of $M$ is $\{|G|=1\}$ for some $G$.
@@
We now show that the leaves of the putative canonical lamination satisfy the necessary curvature bounds; this is a little subtle because the leaves are injectively immersed but not embedded. In the below lemmata, let $r_{*}$ be the minimum of the injectivity radius of $M$ and $\delta\|\mathrm{Riem}_{M}\|_{C^{0}}^{-1/2}$, where $\delta>0$ is a dimensional constant to be determined later. Let $\mathbf{S}^{d-1}$ be the round sphere of dimension d1d-1.
@@
$\mathrm{vol}(\mathbf{B}^{d-1})r^{d-1}\leq\mathrm{vol}(N^{\prime})\leq 2\mathrm{vol}(\mathbf{S}^{d-1})r^{d-1}.$ (4.1)
@@
Let $N^{\prime}$ be a component of $N\cap B(x,r)$, so that $N^{\prime}$ is embedded (not just injectively immersed). By transversality, $N^{\prime}\cap\partial B(x,r)$ is diffeomorphic to a closed d2d-2-dimensional submanifold of $\mathbf{S}^{d-1}$. Since $H_{d-2}(\mathbf{S}^{d-1},\mathbf{R})=0$, there exists a relatively open set $V\subseteq\partial B(x,r)$ which is bounded by $N\cap\partial B(x,r)$. Because of how we chose $r_{*}$, we may use the $L^{\infty}$ Poincaré lemma, Theorem 2.1, to find a continuous d2d-2-form AA on a neighborhood of the closure of $B(x,r)$, such that $F=\mathop{}\!\mathrm{d}A$. Then
@@
Let $x\in N$ and let $r>0$ be small enough depending on $M$. Then each component $N^{\prime}$ of $N\cap B(x,r)$ is $F$-calibrated, and therefore a stable minimal hypersurface. By (4.1), $\mathrm{vol}(N^{\prime})\lesssim r^{d-1}$. So by [44, pg785, Corollary 1] (see also [13, Chapter 2, §§4-5]),
@@
[…]
@@
    $v$ satisfies a linear elliptic PDE on $V$ [13, Proof of Theorem 7.3].
@@
We claim that $v$ is identically 0. If this is not true, the set $\{v=0\}=\{v=\mathop{}\!\mathrm{d}v=0\}$ is d3d-3-rectifiable [29, Lemma 1.9], but $\dim V=d-1$, so $\{v\neq 0\}$ is connected. So either $v\geq 0$ or $v\leq 0$, and $v$ has a zero; this contradicts the maximum principle.
@@
We must show that a limit of $F$-calibrated hypersurfaces is $F$-calibrated, and to make this precise we shall need the notion of a Vietoris limit superior of a sequence of closed sets, [32, §4.F]. If $(K_{n})$ is a sequence of closed subsets of $M$, then $\limsup_{n\to\infty}K_{n}$ is the set of all $x$ such that for every open set $U\ni x$, there exist infinitely many $n\in\mathbf{N}$ such that $U\cap K_{n}\neq\emptyset$; one easily checks that $\limsup_{n\to\infty}K_{n}$ is closed.
@@
By taking a subsequence, we may assume that there exist $x_{n}\in N_{n}$ such that $x_{n}\to x$. By Lemma 4.4, we may also assume that if $N_{n}\cap N_{m}$ is nonempty then $n=m$. Combining this with the curvature bound (4.2), we obtain the hypotheses of [6, Lemma 2.4]. The conclusion of that lemma is that for every $\delta>0$ there exists $r>0$ only depending on $M$, and normal coordinates $(\xi,\eta)\in\mathbf{R}^{d-1}\times\mathbf{R}$ on $B(x,r)$ based at $x$ such that for every $n$,
@@
where one can use [13, (7.21)] to show that $F$ has the form
@@
So by (4.3), we can first choose $\delta$ small enough depending on $M$, and then choose $r$ small enough depending on $\delta$, so that for every $n$ large enough depending on $r$, the equation $Pu_{n}=0$ is uniformly elliptic. In particular, by the interior Schauder estimate [24, Theorem 6.2], we may choose $\delta$ small enough, depending only on $M$, that there exists a connected, relatively open set $W\subseteq V$ such that:
@@
To prove this, let $N_{n}^{x}:=\{(\xi,u_{n}(x)):\xi\in W\}$. If $\mathrm{diam}(W)$ was chosen small enough (depending only on $M$), then we can use the $L^{\infty}$ Poincaré lemma, Theorem 2.1, to find a continuous d2d-2-form AA on $W$ such that $\mathop{}\!\mathrm{d}A=F$. Since $u_{n}\to u$ in $C^{2}(W)$, we can compute using Stokes’ theorem
@@ 5.1. Ergodic theory of $\lambda_{\rho}$
Let $M$ be a closed oriented Riemannian manifold of dimension $2\leq d\leq 7$. For each oriented lamination $\lambda$ in $M$, let $\mathcal{M}(\lambda)$ be the set of transverse probability measures to sublaminations of $\lambda$. This set inherits the vague topology on the space of Borel probability measures on $\operatorname{supp}\lambda$; in view of (2.3), this topology is the same as the topology on the space of measured laminations (see [6]) restricted to $\mathcal{M}(\lambda)$. It is clear that $\mathcal{M}(\lambda)$ is convex, and one may use the compactness of the space of Borel probability measures on the compact metrizable space $\operatorname{supp}\lambda$ [32, Theorem 17.23] to show that $\mathcal{M}(\lambda)$ is compact. By the Krein-Milman theorem, if $\mathcal{M}(\lambda)$ is nonempty, then so is its set of extreme points, $\mathcal{E}(\mathcal{M}(\lambda))$.
@@
The first claim is an easy modification of the proof of [19, Theorem 4.4], and the second is essentially the proof that every ultrafilter on a finite set is principal. To be more precise, let $S$ be a finite set of ergodic measures, and choose $c_{\mu}\in\mathbf{R}$ such that $\sum_{\mu\in S}c_{\mu}\mu=0$. The measures in $S$ are determined by their values on saturated sets, so if some coefficient $c_{\nu}$ is nonzero, then there exists a saturated set $E$ and a proper subset $T\subset S$ such that:
@@ 5.2. The Arnoux–Levitt lemma
Following an idea of Auer and Bangert [3], we study $\mathcal{M}(\lambda_{\rho})$ using an ergodic-theoretic lemma of Arnoux and Levitt, [2, Proposition 3.1]. We need a more general version of the Arnoux–Levitt lemma, and the original proof is in French, so we include a proof here.
Let us identify transverse measures with positive transverse cocycles (cocycles which act on curves transverse to the lamination and are cooriented with the lamination); this is standard, and we refer to [17, §7.2] for a justification of this identification.
@@
We follow [2, §3, Lemme] which is a similar result when $\lambda$ is a minimal component of a foliation.
@@
    NN is dense in $\operatorname{supp}\lambda\cap U$.
@@
    Suppose that $\kappa_{i}$ is not a foliation with a dense leaf. By the Morgan–Shalen decomposition (Theorem 2.12) and the ergodicity of $\mu_{1}^{i}$, $\kappa_{i}$ is exceptional, and if $i^{\prime}\neq i$ then $\operatorname{supp}\kappa_{i^{\prime}}$ avoids an open set $U_{i}$ containing $\operatorname{supp}\kappa_{i}$. Let $N_{i}$ be any leaf of $\kappa_{i}$.
@@
Suppose that there are $a_{j}\in\mathbf{R}$ such that $\sum_{j}a_{j}[\mu^{i}_{j}]=0$, let $\mu$ be the sum of $a_{j}\mu_{i}^{j}$ over $j$ such that $a_{j}>0$, and let $\nu$ be the sum of $a_{j}\mu_{i}^{j}$ over $j$ such that $a_{j}<0$. Then $\mu]-[\nu]=0$, so by Lemma 5.3 with $b=0$ and $U=M$, $\mu-\nu=0$, hence $\sum_{j}a_{j}\mu^{i}_{j}=0$. By ergodicity, $(\mu_{i}^{j})_{j}$ is linearly independent, so $a_{j}=0$, establishing (0–1).
@@
$\mu_{i}^{1}]:H_{1}(M,\mathbf{Z})\to\mathbf{R}.$
Since $M$ is compact, $H_{1}(M,\mathbf{Z})$ is finitely generated and so we can slightly perturb the value of $\mu_{i}^{1}]$ on the generators to obtain $t_{i}$ with the desired properties.
@@
We have already seen that every extreme point of $\mathcal{M}(\lambda)$ is an ergodic measure and so we just need to show the converse. By the Morgan–Shalen decomposition, Theorem 2.12, we may reduce to the case that $\lambda$ either has only closed leaves, or has no closed leaves. If $\lambda$ has only closed leaves, then every ergodic sublamination is a single leaf, and in particular cannot be written as a convex combination of any other ergodic sublaminations. Otherwise, $\lambda$ has no closed leaves, so by Lemma 5.4, $\mathcal{E}(\mathcal{M}(\lambda))$ is finite, say $\{\mu_{1},\dots,\mu_{m}\}$. For each $i\neq j$, there exists a saturated set $E_{ij}$ such that $\mu_{i}(E_{ij})=1$ but $\mu_{j}(E_{ij})=0$. In particular, if $\mu$ is not an extreme point and we write $\mu=\sum_{i}d_{i}\mu_{i}$, then there exist $i\neq j$ such that $d_{i},d_{j}>0$. Then $d_{i}\leq\mu(E_{ij})\leq 1-d_{j}$, so $\mu$ is not ergodic. ∎
@@ 5.3. The dual flat $\rho^{*}$
    $\lambda]$ has rational direction.
@@
If $\lambda$ is a closed hypersurface $N$, then $\lambda]$ is a rescaling of $[N]$, and $[N]$ is the image of the class of $N$ in $H_{d-1}(M,\mathbf{Z})$.
Conversely, assume that $\lambda]$ has rational direction. Since $\pi_{1}(M)$ is finitely generated, we may rescale $M$ suitably so that $\lambda]$ is a representation $\alpha:\pi_{1}(M)\to\mathbf{Z}$. Since such representations are identified with homotopy classes of maps $M\to\mathbf{S}^{1}$, the Ruelle-Sullivan current $T_{\lambda}$ takes the form $\mathop{}\!\mathrm{d}u$ for some map $u:M\to\mathbf{S}^{1}$. Let $\tilde{u}\in BV_{\mathrm{loc}}(\tilde{M},\mathbf{R})$ be the universal cover of $u$.
@@
To see this, let $\tilde{y}\in\partial\{u>t\}\cap B(\tilde{x},r/2)$. Since $\partial\{\tilde{u}>t\}$ is smooth, its density $\theta$ (in the sense of rectifiable sets) at $\tilde{y}$ is the volume of the unit ball of $\mathbf{R}^{d-1}$. By the monotonicity formula for minimal hypersurfaces [36, Theorem 7.11], there exists $A\geq 0$ which only depends on $M$ such that for any $\rho>0$,
@@
The image of $T$ in $\mathbf{S}^{1}$ is a point, so for any $t,s\in T$, either $t=s$ or $|t-s|\geq 1$. We may assume that there is an infinite increasing sequence $(t_{n})$ in $T$. By the coarea formula [25, Theorem 1.23],
@@
So if $\lambda]$ has rational direction, then every leaf of $\lambda$ is closed. Since $\lambda$ is ergodic, it follows that $\lambda$ is a single closed leaf. ∎
Lemma 5.7.
Let $\mathcal{S}$ be a finite set of disjoint closed hypersurfaces in $M$. If $M\setminus\bigcup\mathcal{S}$ is connected, then the set of homology classes of $\mathcal{S}$ is linearly independent over $\mathbf{R}$.
Proof.
Assume towards contradiction that there is a $d$-chain $\sigma$, $N_{1},\dots,N_{k}\in\mathcal{S}$, and $c_{1},\dots,c_{k}\in\mathbf{R}\setminus\{0\}$, such that $\partial\sigma=\sum_{j\leq k}c_{j}N_{j}$. We identify $\sigma$ with the function on the open set $U:=M\setminus\bigcup\mathcal{S}$ which sends a point $x$ to the multiplicity of $\sigma$ at $x$. Then $\sigma$ is locally constant, but $U$ is connected, so $\sigma$ is actually constant. Therefore $\partial\sigma=0$, so $\{N_{1},\dots,N_{k}\}$ is linearly dependent when viewed as a set of d1d-1-chains. So there exists $2\leq j\leq k$ such that $N_{1}\cap N_{j}\neq\emptyset$, a contradiction. ∎
@@
Let $\Pi:\mathcal{M}(\lambda_{\rho})\to\partial B$ be the map which sends a transverse probability measure to its homology class. Let $m$ be the number of measures on sublaminations which are not closed hypersurfaces. By Lemma 5.6, $\Pi$ sends closed hypersurfaces to rational extreme points of $\rho^{*}$ and all other measures to irrational extreme points. By Corollary 1.3, $\Pi$ is surjective, so the number of irrational extreme points is at most $m$. By Lemma 5.4, $m\leq\max(0,b_{1}-1)$.
We claim that $\rho^{*}$ has only finitely many rational extreme points. If not, there are infinitely many closed leaves $N_{n}$ of $\lambda_{\rho}$ with distinct homology classes $\alpha_{n}\in\mathcal{E}(\rho^{*})$. Then $\{\alpha_{1},\dots,\alpha_{b_{1}+1}\}$ is linearly dependent, so by Lemma 5.7, $U:=M\setminus\bigcup_{n\leq b_{1}+1}N_{n}$ must be disconnected. So every noncompact leaf $N$ of $\lambda_{\rho}$ must miss a connected component of the open set $U$; it follows that $N$ is not dense in $M$. In particular, $\lambda_{\rho}$ is not a foliation with a dense leaf.
Let $\mu\in\mathcal{M}(\lambda_{\rho})$ be defined by $\mu(N_{n})=2^{-n}$ (so $\mu(M\setminus\bigcup_{n}N_{n})=0$). Then $\operatorname{supp}\mu$ is a sublamination $\kappa$ of $\lambda_{\rho}$ in which $\bigcup_{n}N_{n}$ is dense, and $\kappa$ is not a foliation with a dense leaf. So by the Morgan–Shalen decomposition (Theorem 2.12), $\kappa$ is the disjoint union of finitely many parallel families of closed leaves. By the pigeonhole principle, there exist $m<n$ such that $N_{m}$ and $N_{n}$ are in the same parallel family. Therefore $\alpha_{m}=\alpha_{n}$, a contradiction; this establishes the claim.
Therefore $\rho^{*}$ is a closed convex set with only finitely many extreme points. So $\rho^{*}$ is a convex polytope. ∎